I am getting stuck and can not figure out what i am doing wrong
the problem is (y-4) - (y+7) = 8y
our teacher is teaching us distributing and with this answer i would just use 1
so i got y - 4 -y -7 + 8y
then i combined like terms
the y's cancel out so now i am left with -11+8y
and this is where i get stuck

Answer by bucky(2189)

I am getting stuck and can not figure out what i am doing wrong
the problem is (y-4) - (y+7) = 8y
.
our teacher is teaching us distributing and with this answer i would just use 1
so i got y - 4 -y -7 + 8y <-- Mistake here. The + sign by the 8y should be an "=". You did the rest correctly in removing the parentheses.
So you should have:
.
y - 4 -y -7 = 8y
.
then i combined like terms. <-- this is the correct thing to do.
the y's cancel out (that's right) so now i am left with -11+8y <--- change + to =
.
So you are left with:
.
-11 = 8y
.
You can reverse the sides (called transposing) to get the equation in a little more conventional form with the unknown (the y term) on the left side and everything else (in this case - 11) on the right side:
.
8y = -11
.
You can now solve for y by dividing both sides of this equation by 8. (You have to divide BOTH sides because whatever you do to one side of the equation you have to do the same thing to the other side to keep things equal).
.
When you divide 8y by 8 you get just y.
.
and when you divide -11 by 8 you get the improper fraction -11/8.
.
and that's the answer you get: y = -11/8
.
You did a good job, but probably changing the = sign to a + sign was a little confusing. You know what you are doing, but sometimes it gets confusing. It will get much easier with more practice. Keep up the good work.
.
Hope this helps.